OpenStudy (anonymous):

help Find the distance between (-1, 4) and (3, 1)

OpenStudy (anonymous):

do you know the distance formula??

OpenStudy (anonymous):

y2-y2 x2-x1?

OpenStudy (anonymous):

\[\sqrt{(y2-y1)^{2}+(x2-x1)^{2}}\]

OpenStudy (anonymous):

can u do it 4 me n i'll give u medaLS on every question u do for me :)

OpenStudy (anonymous):

im not going to do your work for you. you need to learn how to do it. I can't do your test for you. you need to do it now so you know it for later. And i don't really care much for medals. I care to help people. Not to give them the answers

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1-4 ^2 3-1^2 9 4?

OpenStudy (anonymous):

@etemplin

OpenStudy (anonymous):

careful x2-x1 3-(-1) 3+1 4 squared = 16 9+16=25 square root = 5 answer = 5

OpenStudy (anonymous):

was the other one right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and then u add them?

OpenStudy (anonymous):

i did, 9+16=25 square root 25=5 answer = 5

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

wait wer did 16 come frm?

OpenStudy (anonymous):

@etemplin

OpenStudy (anonymous):

x2-x1 3-(-1) 3+1 4 squared = 16

OpenStudy (anonymous):

and the other one? bc tht looks the same as the other one? @etemplin

OpenStudy (anonymous):

the other one is 9.

OpenStudy (anonymous):

how did u get tht 9? like i did?

OpenStudy (anonymous):

yes